OpenStudy (anonymous):

8sqrt{6}-4sqrt{6}+3sqrt{6} 8sqrt{6}-4sqrt{6}+3sqrt{6} @Mathematics

OpenStudy (anonymous):

7sqrt(6)

OpenStudy (anonymous):

\[8\sqrt{6}-4\sqrt{6}+3\sqrt{6}=7\sqrt{16}???\]

OpenStudy (anonymous):

opps 6 i mean

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

u got it right @ tena

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

like this only try to solve as much as possible

OpenStudy (anonymous):

heck that one was easy lol the rest is like i hve noooooo idea

OpenStudy (anonymous):

\[7\sqrt{18}+8\sqrt{50}=7\sqrt{9\times2}+8\sqrt{2\times5\times5}\] thats as far as ive gotten